OpenStudy (anonymous):

Find each absolute value. l 4 + 2i l

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

forget negative numbers, forget \(i\) just compute \(\sqrt{a^2+b^2}\) with \[a=4,b=2\]

OpenStudy (anonymous):

what about the imaginary numbers though?

OpenStudy (anonymous):

what about them?

OpenStudy (anonymous):

like i said, ignore all minus signs and the \(i\) the absolute value is a positive real number, the distance that your complex number is from the origin by pythagoras, it is \(\sqrt{a^2+b^2}\) i can draw a picture if you like

OpenStudy (anonymous):

No it's fine, thankyou!

OpenStudy (anonymous):



OpenStudy (anonymous):

Would the absolute value of l 5 - i l be 5? & ohh okaay I got it :)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no no not 5

OpenStudy (anonymous):

it is \[\sqrt{5^2+1^2}=\sqrt{26}\]

OpenStudy (anonymous):

when i said "forget the \(i\) " i did not mean pretend it is not there, i meant don't include it in your calculations this is the same as \(5-1i\) so you use 5 and 1

OpenStudy (anonymous):

OK Thankyou! :)

OpenStudy (anonymous):

yw